OpenStudy (anonymous):

solve for all values of x \[9/25^{x}=5/3\]

OpenStudy (anonymous):

9/25^x=5/3 27/5 = 25^x ln(27/5) = xln25 3ln3 - ln5 = x(2ln5) x= (3ln3 - ln5)/2ln5